# Strong geometry: knots

### v1

@@ Abstract.
In this paper, we introduce the notion of strong geometry, a structure composed by both the chirotope of the oriented matroid associated to a set of points $X$ in the $d$-dimensional space and a new chirotope induced by the hyperplanes spanned by $X$. With this on hand, we answer positively a strong geometry version of a question due to M. Las Vergnas about reconstructing polygonal knots via oriented matroids.
@@ 1. Introduction
Many problems in discrete and convex geometry are based on finite sets of points in the Euclidean space. The combinatorial properties of the underlying point set, rather than its metric properties, are enough to solve a large number of such problems. Indeed, oriented matroids (initially known as combinatorial geometries) are natural combinatorial objects to study questions concerning the relatively positions of a set of points in the space (not taking into account distances) [1] . For instance, the intersection properties of the line segments spanned by a set of points in the plane turn out to be of great importance. It is well-known that this intersecting properties are exactly reflected by the order type of a point set, introduced in [4]. This geometric condition is well codified combinatorially by the corresponding oriented matroid in terms of the notion of chirotopes (or its signed circuits). Although oriented matroids give information about such intersections, it does not tell anything about the relative positions of such intersections with respect to the set of spanned lines. The knowledge of the latter might play a key role in the study of certain geometric problems. For instance, Figure 1 illustrates two hexagons having the same affine oriented matroid induced by their vertices (they both have the same set of circuits or equivalently set of minimal Radon partitions). However, in the hexagon on the left the lines joining opposite vertices intersect at point $p$ while they do not in the right one.
@@
It turns out that $p$ is a 00-transversal to the convex hull of the 4-sets (that is, $p$ intersects all 4-sets formed by the vertices). It can be easily checked that the hexagon on the right does not admit a 00-transversal. In fact, such discrete transversal is not necessarily an invariant of the order type, see [3].
This is a typical situation in which the knowledge of the position of the intersection of two lines, with respect with the other spanned lines, is helpful (in this case for a transversality-type problem).
In this paper, we introduce the notion of strong geometry. This structure is composed by not only the chirotope of the oriented matroid associated to a set of points $X$ in $\mathbb{R}^{d},d\geq 2$ but also by a new chirotope induced by the hyperplanes spanned by $X$. As we will see, in dimension 2, strong geometries encode nicely the combinatorics of the cells of the arrangement of the spanned lines.
After introducing and discussing properties of the notion of strong geometry, we shall present a first topological application in connection with geometric knots. A knot is an embeddings of $S^{1}$ into $\mathbb{R}^{3}$ up to isotopy. In order to avoid pathologies, it is common to study tame knots, in particular, polygonal knots, that is, closed, piecewise linear loops with no self-intersections consisting of $n$ segments in $\mathbb{R}^{3}$. Not much is known about the interplay between the classic setting and geometric knots, which are polygonal knots with a fixed number of segments of variable length. Geometric knots are more rigid than smooth knots and they better suited for describing macromolecules (such as DNA) in polymer chemistry and molecular biology, but also far more resistant to investigation.
@@
Let $X=(x_{0},\dots,x_{n-1})$ be a $n$-tuple of points in $\mathbb{R}^{3}$ in general position. Is it true that $K_{X}$ only depends on the affine oriented matroid induced by $x_{0},\dots,x_{n-1}$
We are able to give a positive answer to Las Vergnas’ question in the case of strong geometries.
@@ 2. Oriented matroid preliminaries
Let us recall that a set $P=\{p_{1},\dots,p_{n}\}$ of $n\geq d+2$ points in $\mathbb{R}^{d}$ always admits a Radon partition, that is, a partition $P=A\sqcup B$ such that $conv(\{p_{i}\ |\ p_{i}\in A\})\cap conv(\{p_{i}\ |\ p_{i}\in B\})\neq\emptyset$. It is known that a circuit $C=C^{+}\cup C^{-}$ of an affine oriented matroid associated to a set of points $\{y_{1},\dots,y_{n}\}$ corresponds to a minimal Radon partition, that is, $conv(\{y_{i}\ |\ i\in C^{+}\})\cap conv(\{y_{i}\ |\ i\in C^{-}\})\neq\emptyset$. We have that a Radon partition of $\{x_{1},\dots,x_{d+2}\}$ admits both $x_{d+1}$ and $x_{d+2}$ in either $A$ or $B$ if and only if $\chi(x_{1},\dots,x_{d},x_{d+1})=-\chi(x_{1},\dots,x_{d},x_{d+2})$. Therefore, Radon’s partitions can be recovered from the chirotope.
3. Strong geometries
We shall first introduce and discuss the notion of strong geometries when $d=2$ for a better comprehension of this notion. We then treat strong geometries for general $d\geq 3$.
@@ 3.1. Planar case
Let $X$ be a set of points in $\mathbb{R}^{2}$. As mentioned above, strong geometries intend to capture more geometric information than ordinary oriented matroids, specifically, the relative positions of triples of lines generated by points in $X$. Instead of consider the rank 3 affine oriented matroid associated to $X$, we may thought the affine geometry of dimension 2 as a linear geometry in dimension 3. This setting happen to be more convenient, for our propose. Since we are only interested in computing signs of multi-linear expressions then we do not worry about the norm of the vectors in $\mathbb{R}^{3}$. The points in $X$ will be associated to vectors living in $\mathbb{R}^{3}/(\mathbb{R_{+}})$ which in turn will be thought of as $\mathbb{S}^{d}\cup\{0\}$ for a nicer visualisation.
Let $x$ and $y$ be two points in the affine space $\mathbb{R}^{2}\subset\mathbb{R}^{3}$. Let $l_{x,y}$ be the equator arising from the intersection of the (oriented) plane $h_{x,y}$ spanned by $x$ and $y$ and $\mathbb{S}^{3}$. Let $\mathbf{x}$ and $\mathbf{y}$ be the vectors on $l_{x,y}$ arising from intersection of $\mathbb{S}^{2}$ and the line segments $[0,x]$ and $[0,y]$ respectively. Equator $l_{x,y}$ is oriented from $\mathbf{x}$ to $\mathbf{y}$ if $(\mathbf{x},\mathbf{y})$ is a positive basis, see Figure 2.
@@
The hyperplane $h_{x,y}$ is parametrized by the normal positive (norm 1) vector given by the product $\mathbf{x}\times\mathbf{y}$. We may write $\wedge$ the image of $\times$ in $\mathbb{R}^{3}/(\mathbb{R_{+}})$. In the degenerate case, we consider the 0 vector to be the equator between two antipodal (equal) points, see Figure 3.
@@
Let $l$ be an oriented equator of $\mathbb{S}^{3}$ with positive normal vector $v_{l}$ and let $l^{+}$ (resp. $l^{-}$) be the positive (resp. negative) half-sphere delimited by $l$. Let $\mathbf{p}$ be a vector in $\mathbb{S}^{2}$. Then,
@@
Let $x,y$ and $z$ be points in the affine space. One may determine whether $z$ is in the positive or negative half-space delimited by the oriented line from $x$ to $y$ by considering $\mathsf{sign}(\det(x,y,z))$. In a similar manner, this can be determined by considering the vectors $\mathbf{x},\mathbf{y},\mathbf{z}\in\mathbb{S}^{2}$. Indeed, by Remark 1, $\mathsf{sign}(\langle\mathbf{x}\times\mathbf{y},\mathbf{z}\rangle)$ determines whether $\mathbf{z}$ is in the positive or negative half-sphere delimited by $l_{x,y}$ oriented from $x$ to $y$. The latter is simply due to the elementary property of the determinant
(1) $\mathsf{sign}(\det(x,y,z))=\mathsf{sign}(\langle\mathbf{x}\times\mathbf{y},\mathbf{z}\rangle).$
Dealing with equators in the sphere makes things easier. In affine geometry there are numerous ways that three oriented lines may met, see Figure 4 while generic triplets of oriented equators can only be of two combinatorial types, see Figure 5.
@@
Let $l_{1}$ and $l_{2}$ be two equators and let $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ be the normal vectors of the corresponding hyperplanes supporting $l_{1}$ and $l_{2}$ respectively. If $l_{1}$ and $l_{2}$ are colinear, i.e. equal or opposite, then the sign associated to any triple $(l_{1},l_{2},l)$ is always 00. Otherwise, they meet at two antipodal points $p$ and $-p$, in this case, $\mathbf{v}_{1}\wedge\mathbf{v}_{2}$ is one of these points, we call it positive intersection, see Figure 6.
@@
We thus have that a third equator $l_{3}$ admits either $p$ or $-p$ on its positive half-sphere (the other in the negative half-sphere). Analogously to the dual case above (1), we define the sign associated to a triple of equators $l_{1},l_{2}$ and $l_{3}$ as
(2) $\mathsf{sign}(l_{1},l_{2},l_{3}):=\mathsf{sign}(\langle\mathbf{v}_{1}\times\mathbf{v}_{2}\rangle,\mathbf{v}_{3}\rangle).$
@@ 3.2. General setting
Let $d\geq 2$ and let $X=(\mathbf{x}_{1},\dots,\mathbf{x}_{d})$ be a $d$-tuple of vectors in $\mathbb{R}^{d+1}$ (or in $\mathbb{S}^{d}$). If the family $(\mathbf{x}_{1},\dots,\mathbf{x}_{d})$ is independent then we define $\boldsymbol{\alpha}(X)$ to be the (norm 1) $(d+1)$-dimensional vector orthogonal to all $\mathbf{x}_{i}$’s (i.e., orthogonal to the hyperplane $h(X)$ spanned by the $\mathbf{x}_{i}$’s) and such that $\det(\mathbf{x}_{1},\dots,\mathbf{x}_{d},\boldsymbol{\alpha}(X))>0$. We call $\boldsymbol{\alpha}(X)$ an hyperplane-vector. If $(\mathbf{x}_{1},\dots,\mathbf{x}_{d}$ is dependent then we set $\boldsymbol{\alpha}(X)=\mathbf{0}$ (in this case the hyperplane vector is degenerate).
@@
We define the wedge oriented matroid, denoted by $M_{\mathsf{Wedge}}(X)$, as the oriented matroid $M_{\mathsf{Lin}}(\Lambda(X))$, that is, the $(d+1)$-rank oriented matroid with base set $[n]^{d}$ and with chirotope function $\chi_{\Lambda}$ verifying
@@
Let $X_{1},\dots,X_{d},Y$ be $(d+1)$ $d$-tuples of vectors in $\mathbb{R}^{d}$ and let $Y=(\mathbf{y}_{1},\dots,\mathbf{y}_{d})$. Then,
@@
the desired equality follows. ∎
Figure 7 gives a visual explanation of Proposition 1 in the affine setting.
Figure 7.
We notice that Proposition 1 implies that two $n$-tuples with the same strong geometry have either equal or opposite chirotopes. Let us extend Proposition 1 to affine space in $\mathbb{R}^{4}$. For, we first introduce the witness oriented matroid which will play a central role in the proof of our main result. Let $\omega,x_{1},\dots,x_{n}$ be a configuration of points in the affine space $\mathbb{R}^{d}$ where the point $\omega$, called the witness, will play a special role.
The witness oriented matroid from $\omega$, denoted by $M_{\mathsf{Wit}_{\omega}}(X)$, is the linear oriented matroid $M_{\mathsf{Lin}}(W_{\omega}(X))$ where $W_{\omega}(X)=\{x_{1}-\omega,\cdots,x_{n}-\omega$}. Notice that $M_{\mathsf{Wit}_{\omega}}(X)$ is of rank $d$ since the vectors $x_{i}-\omega$ are considered in $\mathbb{R}^{d}$ instead of vectors in $\mathbb{R}^{d+1}$, see Figure 8.
@@
Let $n\geq 1$ be an integer and let $X=(\omega=x_{0},x_{1},\dots,x_{n})$ be a tuple of points in $\mathbb{R}^{3}$. Let $i_{0}=0$, $i_{1},\dots,i_{9}\in[n]^{9}$ and let $I_{1}=(i_{0},i_{1},i_{2})$, $I_{2}=(i_{0},i_{3},i_{4})$, $I_{3}=(i_{0},i_{5},i_{6})$ and $J=(i_{7},i_{8},i_{9})$. Let $\mathbf{x}_{a_{1}}=(x_{i_{1}}-\omega)\wedge(x_{i_{2}}-\omega),\mathbf{x}_{a_{2}}=(x_{i_{3}}-\omega)\wedge(x_{i_{4}}-\omega)$ and $\mathbf{x}_{a_{3}}=(x_{i_{5}}-\omega)\wedge(x_{i_{6}}-\omega)$. Then,
$\chi_{\Lambda}(I_{1},I_{2},I_{3},J)=+\chi_{\omega}(a_{1},a_{2},a_{3})\chi(i_{0},i_{7},i_{8},i_{9})$
where $\chi_{\Lambda},\chi$ and $\chi_{\omega}$ are the chirotopes of the rank 4 matroids $M_{\mathsf{Wedge}}(\mathsf{in}(X))$ and $M_{\mathsf{Lin}}(\mathsf{in}(X))=M_{\mathsf{Aff}}(X)$ and the rank 3 matroid $M_{\mathsf{Wit}_{\omega}}(X)$ respectively.

### v2

@@ Abstract.
In this paper, we introduce the notion of strong geometry, a structure composed by both the chirotope of a set of points $X$ in the $d$-dimensional space and the wedge chirotope which is the specific adjoint chirotope induced by the hyperplanes spanned by $X$. We present various properties relating these two chirotopes, for instance, by introducing the witness chirotope, we are able to give a formula expressing the wedge chirotope in terms of the usual chirotope. With this on hand, we answer positively a strong geometry version of a question due to M. Las Vergnas about reconstructing polygonal knots via chirotopes.
@@ 1. Introduction
Many problems in discrete and convex geometry are based on finite sets of points in the Euclidean space. The combinatorial properties of the underlying point set, rather than its metric properties, are enough to solve a large number of such problems. The codification of the ‘combinatorial properties’ of a point set usually refers to a natural associated notion called chirotope (initially known as combinatorial geometry, with other equivalent names including order type [12] and, in our case, oriented matroid). The chirotope captures the relative positions of small tuples of points without taking distances into account [4]. In the affine plane, one such piece of information we can formulate in terms of chirotopes is the intersection properties of the segments spanned by the initial points. However, even though the chirotope tells us which segments intersect, it does not give any information about the relative positions of those intersection (say with respect to the spanned lines). The knowledge of the latter might play a key role in the study of certain geometric problems. For instance, Figure 1 illustrates two hexagons having the same chirotope induced by their vertices (they both have the same set of circuits or equivalently set of minimal Radon partitions). However, in the hexagon on the left the lines joining opposite vertices intersect at point $p$ while they do not on the right one.
@@
It turns out that $p$ is a 0-transversal to the convex hull of the 4-sets (that is, $p$ intersects all 4-sets formed by the vertices). It can be easily checked that the hexagon on the right does not admit a 0-transversal. In fact, such discrete transversal is not necessarily an invariant of the order type, see [6]. This is a typical situation in which the knowledge of the position of the intersection of two lines, with respect with the other spanned lines, is helpful (in this case for a transversality-type problem).
Let $X$ be a vector configuration of $\mathbb{R}^{d},d\geq 3$. We introduce the notion of strong geometry $\mathsf{SGeom}_{\mathsf{Lin}}(X)$ associated to $X$. This structure is composed of both the usual linear $M_{\mathsf{Lin}}(X)$ and the wedge $M_{\mathsf{Wedge}}(X)$ chirotopes associated to $X$. The wedge chirotope is a specific adjoint chirotope of $M_{\mathsf{Lin}}(X)$ associated to $X$ arising from the spanning vector configuration consisting of exactly one non-zero vector orthogonal to each hyperplane in $\mathbb{R}^{d}$ that is spanned by a subset of $X$, see [8, Section 3] (this canonical construction will be treated in detail below). This adjoint lives in the same dimension, but typically it has many more vectors than the original configuration. It is known that every rank 3 oriented matroid has an adjoint [11, 9, 13] but not every oriented matroid has one, for instance, the Vámos matroid has no oriented adjoint [3]. Although a realizable oriented matroid $M$ always admits an adjoint, the latter is not uniquely determined by $M$, it depends on the particular vector configuration representing $M$ (see below). We refer the reader to [1, 7] and [4, Section 5.3] for further discussions on adjoints for general oriented matroids.
Strong geometries give helpful information, for instance, they encode nicely the combinatorics of the cells of the arrangement of the hyperplanes spanned by a set of points. In Section 6, we discuss different contexts closely connected to strong geometries.
In the next section, we treat in detail wedge chirotopes. In rank 333, that is, in the affine plane, we give a formula expressing the wedge chirotope in terms of the usual chirotope for some specific given configuration (see Eq. (3) and Proposition 1). Our approach allows us to get a ‘dimension reduction’ formula for any dimension. We will show that the chirotope of $M_{\mathsf{Wedge}}(X)$ determines the chirotope of $M_{\mathsf{Aff}}(X)$ up to orientation depending on the parity of the dimension (see Proposition 2). In order to show the latter, we introduce the witness chirotopes which play a central role for our purpose.
We then focus our attention on the study of a first topological application of strong geometries in connection with geometric knots. A knot is an embeddings of $S^{1}$ into $\mathbb{R}^{3}$ up to ambiant isotopy. In order to avoid pathologies, it is common to study tame knots, in particular, polygonal knots, that is, closed, piecewise linear loops in $\mathbb{R}^{3}$ with no self-intersections. Not much is known about the interplay between the classic setting and geometric knots, which are polygonal knots with a fixed number of segments of variable length. Geometric knots are more rigid than smooth knots and are better suited to describe macromolecules (such as DNA) in polymer chemistry and molecular biology, but also far more resistant to investigation.
@@
Let $X=(x_{0},\dots,x_{n-1})$ be a $n$-tuple of points in $\mathbb{R}^{3}$ in general position. Is it true that $K_{X}$ only depends on the chirotope induced by $x_{0},\dots,x_{n-1}$
By considering the affine version of strong geometries arising from a set of points $X$ in the space, we are able to give a positive answer to a strong geometry version of Las Vergnas’ question. More conveniently, the result is stated in the following equivalent way.
@@
Since the chirotope of $M_{\mathsf{Wedge}}(X)$ determines the chirotope of $M_{\mathsf{Lin}}(X)$ in dimension 4 (see Corollary 3) then the following result is a straightforward consequence of Theorem 1.
Corollary 1.
Let $X=(x_{0},\dots,x_{n-1})$ and $X^{\prime}=(x^{\prime}_{0},\dots,x^{\prime}_{n-1})$ be two $n$-tuples of points in $\mathbb{R}^{3}$ in general position. If $M_{\mathsf{Wedge}}(X)$ is isomorphic to $M_{\mathsf{Wedge}}(X^{\prime})$ then $K_{X}$ is isotopic $K_{X^{\prime}}$.
In order to obtain Theorem 1, we prove that knotoids are determined by their Gauss diagram (see Lemma 3). The arguments for the latter are extended to show that graphoids are also determined by their Gauss diagram (see Lemma 4). This allows us to show that linear spatial graphs are determined by their corresponding strong geometries as well (see Theorem 2).
@@ 2. Oriented matroid preliminaries
Let us recall that a set $P=\{p_{1},\dots,p_{n}\}$ of $n\geq d+2$ points in $\mathbb{R}^{d}$ always admits a Radon partition, that is, a partition $P=A\sqcup B$ such that $conv(\{p_{i}\ |\ p_{i}\in A\})\cap conv(\{p_{i}\ |\ p_{i}\in B\})\neq\emptyset$. It is known that a circuit $C=C^{+}\cup C^{-}$ of an affine oriented matroid associated to a set of points $\{y_{1},\dots,y_{n}\}$ corresponds to a minimal Radon partition, that is, $conv(\{y_{i}\ |\ i\in C^{+}\})\cap conv(\{y_{i}\ |\ i\in C^{-}\})\neq\emptyset$. If $(x_{1},\dots,x_{d+1})$ is generic then a Radon partition of $\{x_{1},\dots,x_{d+2}\}$ admits both $x_{d+1}$ and $x_{d+2}$ in either $A$ or $B$ if and only if $\chi(x_{1},\dots,x_{d},x_{d+1})=-\chi(x_{1},\dots,x_{d},x_{d+2})$. Therefore, Radon’s partitions can be recovered from the chirotope.
3. Wedge matroids
We shall first discuss the construction of wedge oriented matroids for $d=2$ for a better comprehension. We then treat the general case $d\geq 3$.
@@ 3.1. Planar case
Let $X$ be a set of points in $\mathbb{R}^{2}$. As mentioned above, strong geometries intend to capture more geometric information than ordinary chirotopes, specifically, the relative positions of triples of lines generated by points in $X$. As explained above, it will be more convenient to consider $X$ as included in $\mathbb{R}^{3}$. Since we are only interested in computing signs of multi-linear expressions then we do not worry about the norm of the vectors in $\mathbb{R}^{3}$. Therefore, the points in $X$ will be associated to vectors living in $\mathbb{R}^{3}/(\mathbb{R_{+}})$ which is $\mathbb{S}^{d}\cup\{0\}$.
[…]
@@
If we write $\wedge$ for the image of $\times$ in $\mathbb{R}^{3}/(\mathbb{R_{+}})$ then the equator $l_{x,y}$ is parametrized by its normal positive vector given by the product $\mathbf{x}\wedge\mathbf{y}$. In the degenerate case, we consider the 0 vector to be the equator between two equal or antipodal points, see Figure 3.
@@
Let $l$ be an oriented equator of $\mathbb{S}^{2}$ with positive normal vector $\mathbf{v}_{l}$ and let $l^{+}$ (resp. $l^{-}$) be the positive (resp. negative) half-sphere delimited by $l$. Let $\mathbf{p}$ be a vector in $\mathbb{S}^{2}$. Then,
@@
Let $x,y$ and $z$ be points in the affine space. We commonly interpret $\Delta\left(\begin{pmatrix}1&1&1\\ x&y&z\end{pmatrix}\right)$ as indicating whether $z$ lies on the positive or negative half-space delimited by the line from $x$ to $y$. Similarly, this sign is also given by either
(1) $\Delta(\mathbf{x},\mathbf{y},\mathbf{z})=\mathsf{sign}(\langle\mathbf{x}\wedge\mathbf{y},\mathbf{z}\rangle),$
or the sign of the half-sphere delimited by the equator from $\mathbf{x}$ to $\mathbf{y}$ that contains $\mathbf{z}$
These observations serve as a guideline for a natural definition of the combinatorics of a tuple of oriented equators. Indeed, in affine geometry there are numerous ways that three oriented lines may met, see Figure 4 while generic triplets of oriented equators can only be of two combinatorial types, see Figure 5.
@@
[…]
@@
[…]
(2) $\mathsf{sign}(l_{1},l_{2},l_{3}):=\mathsf{sign}(\langle\mathbf{v}_{1}\wedge\mathbf{v}_{2},\mathbf{v}_{3}\rangle).$
@@ 3.2. General setting
[…]
@@
[…]
@@
[…]
@@
the desired equality follows, see Figure 7. ∎
Figure 7. A visual explanation of Proposition 1 in the affine setting.
A consequence of Proposition 1 is the following
Corollary 2.
Let $X$ and $X^{\prime}$ be two $n$-tuples in $\mathbb{R}^{3}$. If $M_{\mathsf{Wedge}}(X)$ is isomorphic to $M_{\mathsf{Wedge}}(X^{\prime})$ then the chirotopes $M_{\mathsf{Lin}}(X)$ and $M_{\mathsf{Lin}}(X^{\prime})$ are either equal or opposite.
Proof.
[…]
We notice that, conversely, two opposite configurations have the same wedge chirotope.
We shall extend Proposition 1 to affine space $\mathbb{R}^{3}$. To this end, we introduce the witness chirotope but, before doing so, let us first briefly discuss our motivation behind this new structure.
The main issue is that, instead of studying knots through their diagrams (orthogonal projections in the plane), we rather consider a projection of the knot radially to a sphere, say with center $\omega$. Therefore, given $x_{1},\dots,x_{n}\in\mathbb{R}^{3}$, we are interested in the linear geometry of the $x_{1}-\omega,\dots,x_{n}-\omega\in\mathbb{R}^{3}$ (notice that the rank decreases by one). In this setting, the most relevant piece of information in the knot sphere diagram will be given by the strong geometry of such configuration (that is, a strong geometry witnessed by $\omega$). As we will see in Proposition 2, the latter is induced by the strong geometry of the base configuration since affine oriented hyperplanes $\alpha(\mathbf{\omega},\mathbf{x}_{1},\mathbf{y}_{1}),\alpha(\mathbf{\omega},\mathbf{x}_{2},\mathbf{y}_{2}),\alpha(\mathbf{\omega},\mathbf{x}_{3},\mathbf{y}_{3})$ meet either positively or negatively at $\omega$ depending on whether $\mathbf{\omega}$ sees the oriented lines $(\mathbf{x}_{1},\mathbf{y}_{1}),(\mathbf{x}_{2},\mathbf{y}_{2}),(\mathbf{x}_{3},\mathbf{y}_{3})$ as a positive or negative triple. Notice that Proposition 1 is the rewording of the preceding remark in lower rank, although the strong geometry is hidden by the fact that, in rank 222, it is the same as the usual chirotope.
The above discussion naturally leads us to define the witness chirotope as follows. Let $\omega,x_{1},\dots,x_{n}$ be a configuration of points in affine space $\mathbb{R}^{d}$, we note $\chi_{\omega}$ the linear chirotope of the vector configuration $(x_{i}-\omega)_{i\in[n]}$. Of course, since
$\left|\begin{matrix}1&1&\dots&1\\ \omega&x_{i_{1}}&\dots&x_{i_{d}}\end{matrix}\right|=\left|\begin{matrix}1&0&\dots&0\\ \omega&x_{i_{1}}-\omega&\dots&x_{i_{n}}-\omega\end{matrix}\right|=|x_{i_{1}}-\omega,\dots,x_{i_{n}}-\omega|,$
then $\chi_{\omega}(\cdot,\dots,\cdot)=\chi(\omega,\cdot,\dots,\cdot)$, see figure 8.
@@
Since we want to study ‘witnessed strong geometry’, it is also natural to define $\chi_{\omega,\wedge}$ on $(x_{i})\in(\mathbb{R}^{d})^{n}$ as the rank $d$, linear wedge chirotope of $(x_{i}-\omega)_{i\in[n]}$
In order to present a clearer and more proper proofs, it is convenient to have a second approach to define a ‘witness chirotope’ that starts and ends in a linear space and, of course, coinciding with the above definition.
Since $x\mapsto x-\omega$ simply acts on $\{1\}\times\mathbb{R}^{d}$ as the projection on $\{0\}\times\mathbb{R}^{d}$ with kernel $<\omega>$, the most obvious replacement is the orthogonal projection $\pi_{\omega^{\perp}}$ on $\omega^{\perp}$. In order to be consistent with the previous definition, we orient $\omega^{\perp}$ such that positive bases of $\omega^{\perp}$ are the $(b_{1},\dots,b_{d-1})$ such that $(\omega,b_{1},\dots,b_{d-1})$ is a positive basis of $\mathbb{R}^{d}$. We thus have that the linear equivalent of the above definition can be obtained by setting the witnessed strong geometry of $X$ witnessed by $\omega$ as the strong geometry of $\pi_{\omega^{\perp}}(X)$ (in $\omega^{\perp}\simeq\mathbb{R}^{d-1}$).
Although both definitions seem to be the most natural ones, it is not immediate to convince one-self that they coincide. Indeed, if all the $\omega,x_{i}$ live in $\{1\}\times\mathbb{R}^{3}$, then the map $x\mapsto x-\omega$ acts on the $x_{i}$ as the projection on $\{0\}\times\mathbb{R}^{3}$ with kernel $\omega$, which is not necessarily the orthogonal projection (unless $\omega=e_{1}$ of course)
Let us check that both projections of the point configurations do have the same strong geometry.
Lemma 2.
Let $H_{1},H_{2}$ be two hyperplanes of $\mathbb{R}^{d}$ not containing $\omega\neq 0$, and let $\pi_{1},\pi_{2}$ the projections with kernel $<\omega>$ and respective images $H_{1},H_{2}$. Suppose further that the $H_{i}$ are oriented such that a positive base $(b_{1},\dots,b_{d-1})$ of $H_{i}$ induces a positive base $(\omega,b_{1},\dots,b_{d-1})$ of $\mathbb{R}^{d}$. Then for any $X\in(\mathbb{R}^{d})^{n}$, $\pi_{1}(X)$ and $\pi_{2}(X)$ have the same strong geometry.
Proof.
We first notice that the $\pi_{i}(X)$ only differ by a positive isomorphism. Indeed the composition $H_{1}\hookrightarrow\mathbb{R}^{d}\rightarrow H_{2}$ of the canonical inclusion and the projection $\pi_{2}$ is a linear map between vector spaces of same dimensions, with trivial kernel, that is, an isomorphism. It also sends a basis $(b_{1},\dots,b_{d-1})$ to some $(b_{1}-\lambda_{1}\omega,\dots,b_{d-1}-\lambda_{d-1}\omega)$, and since $\left|\omega,b_{1}-\lambda_{1}\omega,\dots,b_{d-1}-\lambda_{d-1}\omega\right|=\left|\omega,b_{1},\dots,b_{d-1}\right|$, it is a positive isomorphism. We claim that a positive isomorphism preserve the strong geometry.
Firstly, notice that for a family $(h_{i})_{i\in I}$ of hyperplanes and $(v_{i})_{i\in I}$ its family of normal vectors, a vector $v$ is orthogonal to all $v_{i}$ if and only if it belongs to all $h_{i}$. But a $((d-1)\times(d-2))$-tuple $(x_{i,j})$ verifies $\Delta\Big{(}\alpha(x_{1,1},\dots,x_{1,d-2}),\dots,\alpha(x_{d-1,1},\dots,x_{d-1,d-2})\Big{)}=0$ if and only if the $\alpha(x_{i,1},\dots,x_{i,d-2})$ are linearly dependent, or equivalently, if there is a non-zero vector orthogonal to all of them. Therefore, $\Delta\Big{(}\alpha(x_{1,1},\dots,x_{1,d-2}),\dots,\alpha(x_{d-1,1},\dots,x_{d-1,d-2})\Big{)}=0$ if and only if $\mbox{dim}\left(\bigcap_{i}<x_{i,1},\dots,x_{i,d-2}>\right)\neq 0$. Now acting on the $x_{i,j}$ by isomorphism does not change this dimension, therefore
$\Delta\Big{(}\alpha(x_{1,1},\dots,x_{1,d-2}),\dots,\alpha(x_{d-1,1},\dots,x_{d-1,d-2})\Big{)}=0$
if and only if
$\Delta\Big{(}\alpha(\phi(x_{1,1}),\dots,\phi(x_{1,d-2})),\dots,\alpha(\phi(x_{d-1,1}),\dots,\phi(x_{d-1,d-2}))\Big{)}=0$
for any isomorphism $\phi$.
Take now a positive isomorphism $\phi$. Since the group of positive isomorphisms is path connected, we have a continuous path $\phi_{t}$ of (positive) isomorphisms such that $\phi_{0}=\mbox{id}$, $\phi_{1}=\phi$. If we had, say
$\Delta\Big{(}\alpha(x_{1,1},\dots,x_{1,d-2}),\dots,\alpha(x_{d-1,1},\dots,x_{d-1,d-2})\Big{)}>0$
and
$\Delta\Big{(}\alpha(\phi(x_{1,1}),\dots,\phi(x_{1,d-2})),\dots,\alpha(\phi(x_{d-1,1}),\dots,\phi(x_{d-1,d-2}))\Big{)}<0$
on some $(x_{i,j})$, then there would be some $t\in(0,1)$ such that
$\Delta\Big{(}\alpha(\phi_{t}(x_{1,1}),\dots,\phi_{t}(x_{1,d-2})),\dots,\alpha(\phi_{t}(x_{d-1,1}),\dots,\phi_{t}(x_{d-1,d-2}))\Big{)}<0,$
but, we just saw that this cannot happen. Hence,
$\Delta\Big{(}\alpha(x_{1,1},\dots,x_{1,d-2}),\dots,\alpha(x_{d-1,1},\dots,x_{d-1,d-2})\Big{)}=\Delta\Big{(}\alpha(\phi(x_{1,1}),\dots,\phi(x_{1,d-2})),\dots,\alpha(\phi(x_{d-1,1}),\dots,\phi(x_{d-1,d-2}))\Big{)}.$
Implying that positive isomorphisms preserve the strong geometry, as desired. ∎
We have thus checked that the natural definition of witnessed strong geometry from affine space to linear space coincides with the definition from linear space to linear space. A rigorous formulation of the latter is given by the following generalized version of Proposition 1.
@@
Let $\omega,x_{1,1},\dots,x_{1,d-2},\dots,x_{d-1,d-2},y_{1},\dots,y_{d-1}\in\mathbb{R}^{d}$, with $\omega\neq 0$. Let us denote by $\tilde{\alpha}$ and $\tilde{\Delta}$ the $\alpha$ and $\Delta$ functions associated with the oriented hyperplane $\omega^{\perp}$ Then, we have that
$\Delta\Big{(}\alpha(\omega,x_{1,1},\dots,x_{1,d-2}),\dots,\alpha(\omega,x_{d-1,1},\dots,x_{d-1,d-2}),\alpha(y_{1},\dots,y_{d-1})\Big{)}$
$=(-1)^{d+1}$ $\tilde{\Delta}\Big{(}\tilde{\alpha}(x_{1,1},\dots,x_{1,d-2}),\dots,\tilde{\alpha}(x_{d-1,1},\dots,x_{d-1,d-2})\Big{)}\Delta(\omega,y_{1},\dots,y_{d-1})$